# Enhancing Heterogeneous Multi-Agent Cooperation in Decentralized MARL via GNN-driven Intrinsic Rewards

### 10 Mar 2026 v3

@@ 3.1. Markov Games
where, $\mathcal{V}=\{1,2,\ldots,N\}$ represents the set of all $N$ agents. The state space is denoted by $\mathcal{S}$. The observation space is defined as $\mathcal{O}\equiv\mathcal{O}_{1}\times\mathcal{O}_{2}\times\ldots\times\mathcal{O}_{n}$, where $\mathcal{O}_{i}\subseteq\mathcal{S}$ for each agent $i$ in $\mathcal{V}$. The action space for all agents is denoted by $\mathcal{A}\equiv\mathcal{A}_{1}\times\mathcal{A}_{2}\times\ldots\times\mathcal{A}_{n}$. Each agent $i\in\mathcal{V}$ has an observation instance $o_{i}\in\mathcal{O}_{i}$, which represents a partial view of $\mathcal{S}$. The reward function for agent $i$ is given by $\mathcal{R}_{i}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathcal{R}$. The stochastic state transition model is denoted by $\mathcal{T}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$, describing how each agent transitions to the next state, given the current state and the action taken. Finally, $\gamma$ represents the discount factor, where $0\leq\gamma\leq 1$. Additionally, the reward that agent $i$ receives at time step $t$ can be denoted using $r_{i}^{t}=\mathcal{R}_{i}(s^{t},a^{t},s^{t+1})$, where $a^{t}=(a_{1}^{t},a_{2}^{t},\ldots,a_{n}^{t})$ is the joint action taken by all the agents at time step $t$. The goal of each agent is to maximize the total expected discounted return $R_{i}=\sum^{T}_{t=0}\gamma^{t}r_{i}^{t}$ over the course of an episode with horizon $T$.
@@ 3.2. Message-passing Graph Neural Network
In Equation 1, $\psi_{\theta_{i}}$ and $\phi_{\theta_{i}}$ are two MLPs parameterized by $\theta_{i}$ and the aggregation operator $\bigoplus$ sums the $\phi_{\theta_{i}}$ outputs for all the neighbors of agent $i$. Finally, two distinct MLP decoders take the GNN output $h_{i}$ and produce the value $V_{i}(o_{\mathcal{N}_{i}})$ and the action $a_{i}$, distributed according to $a_{i}\sim\pi_{i}(\cdot\mid o_{\mathcal{N}_{i}})$. This formulation of GNN allows us to utilize it for both intrinsic reward calculation and heterogeneous policy learning of HetGPPO, based on local neighborhood information.
@@ 4.1. Algorithm Description
Figure 2. The per-agent dynamics models in Figure 2(a) are used for calculating the intrinsic rewards, which are then combined with the extrinsic reward from the environment, resulting in $r_{total_{i}}$ for each agent $i$. This combined reward is passed to the HetGPPO policy learning module in Figure 2(b) for heterogeneous policy learning
@@ 5.1. Environments
  • VMAS Flocking: $N$ agents flock around a landmark with random obstacles. Higher velocity and smaller inter-agent distances are rewarded. Collisions are penalized.
  • MPE Simple Spread: $N$ agents aim to occupy $N$ distinct landmarks, receiving rewards for doing so and incurring penalties for collisions.
  • VMAS Reverse Transport: $N$ agents collaboratively push a heavy package toward a random goal within an enclosure, with rewards given upon reaching the goal.
@@
  • VMAS Sampling: $N$ agents start in a grid environment with one-time rewards. Heterogeneous agents use LIDARs with varying ranges and face partial observability.
@@ 5.4. Reward Architecture Evaluation
5.5. Robustness to Increasing No. of Agents

### 12 Mar 2026 v4

@@ 3.1. Markov Games
where, $\mathcal{V}=\{1,2,\ldots,N\}$ represents the set of all NN agents. The state space is denoted by $\mathcal{S}$. The observation space is defined as $\mathcal{O}\equiv\mathcal{O}_{1}\times\mathcal{O}_{2}\times\ldots\times\mathcal{O}_{n}$, where $\mathcal{O}_{i}\subseteq\mathcal{S}$ for each agent $i$ in $\mathcal{V}$. The action space for all agents is denoted by $\mathcal{A}\equiv\mathcal{A}_{1}\times\mathcal{A}_{2}\times\ldots\times\mathcal{A}_{n}$. Each agent $i\in\mathcal{V}$ has an observation instance $o_{i}\in\mathcal{O}_{i}$, which represents a partial view of $\mathcal{S}$. The reward function for agent $i$ is given by $\mathcal{R}_{i}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto\mathcal{R}$. The stochastic state transition model is denoted by $\mathcal{T}\colon\mathcal{S}\times\mathcal{A}\times\mathcal{S}\mapsto[0,1]$, describing how each agent transitions to the next state, given the current state and the action taken. Finally, $\gamma$ represents the discount factor, where $0\leq\gamma\leq 1$. Additionally, the reward that agent $i$ receives at time step $t$ can be denoted using $r_{i}^{t}=\mathcal{R}_{i}(s^{t},a^{t},s^{t+1})$, where $a^{t}=(a_{1}^{t},a_{2}^{t},\ldots,a_{n}^{t})$ is the joint action taken by all the agents at time step $t$. The goal of each agent is to maximize the total expected discounted return $R_{i}=\sum^{T}_{t=0}\gamma^{t}r_{i}^{t}$ over the course of an episode with horizon $T$.
@@ 3.2. Message-passing Graph Neural Network
In Equation 1, $\psi_{\theta_{i}}$ and $\phi_{\theta_{i}}$ are two MLPs parameterized by $\theta_{i}$ and the aggregation operator $\bigoplus$ sums the $\phi_{\theta_{i}}$ outputs for all the neighbors of agent ii. Finally, two distinct MLP decoders take the GNN output $h_{i}$ and produce the value $V_{i}(o_{\mathcal{N}_{i}})$ and the action $a_{i}$, distributed according to $a_{i}\sim\pi_{i}(\cdot\mid o_{\mathcal{N}_{i}})$. This formulation of GNN allows us to utilize it for both intrinsic reward calculation and heterogeneous policy learning of HetGPPO, based on local neighborhood information.
@@ 4.1. Algorithm Description
Figure 2. The per-agent dynamics models in Figure 2(a) are used for calculating the intrinsic rewards, which are then combined with the extrinsic reward from the environment, resulting in $r_{total_{i}}$ for each agent ii. This combined reward is passed to the HetGPPO policy learning module in Figure 2(b) for heterogeneous policy learning
@@ 5.1. Environments
  • VMAS Flocking: NN agents flock around a landmark with random obstacles. Higher velocity and smaller inter-agent distances are rewarded. Collisions are penalized.
  • MPE Simple Spread: NN agents aim to occupy NN distinct landmarks, receiving rewards for doing so and incurring penalties for collisions.
  • VMAS Reverse Transport: NN agents collaboratively push a heavy package toward a random goal within an enclosure, with rewards given upon reaching the goal.
@@
  • VMAS Sampling: NN agents start in a grid environment with one-time rewards. Heterogeneous agents use LIDARs with varying ranges and face partial observability.
@@ 5.4. Reward Architecture Evaluation
$5.$ Robustness to Increasing No. of Agents